The expected return on the market is 17.72 percent, the risk free rate is 4.25 perfent, and tax rate is 25%. Fairfax Paint has 800,000 common shares outstanding that are priced at 35 dollars per share and have an expected return of 16.5% and an expected real return of 14.77%. Last year, Fairfax Paint common stock had a return of 18.3%. The company also had 600,000 shares of preferred stock outstanding that are priced at 20 dollars per share and have an expected return of 11.7% and an expected real return of 10.4%. Last year FP preferred stock had a return of 13.42%. Finally, the company has 10,000 bonds outstanding with a coupon rate of 10.27%, yield to maturity of 4.89%, current yoeld of 9.34%, face value of 1000 dollars, and price of 1100 dollars. What is the weighted average cost of capital for Fairfax Paint? Answer as a rate in decimal format so that 12.34% would be entered as .1234.
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASEC w ENG 06:19 31-07-2020 99+ 29 Х B1023 А. B с D E F G — I FAIRFAX AFTER TAX COST OF DEBT = YTM*(1-t) = 4.89%*(1-0.25) = 3.66